Two investments earn an annual income of $575. One investment earns an annual simple interest rate of 8.5%, and the other investment earns an annual simple interest rate of 6.4%. the amount invested is $8000. How much is invested in each account?
Answer by srinivas123(23)

Two investments earn an annual income of $575. One investment earns an annual simple interest rate of 8.5%, and the other investment earns an annual simple interest rate of 6.4%. the amount invested is $8000. How much is invested in each account?
let the amount given for 8.5 % simple interest be x.
Then it is obvious that the other amount is 8000-x.
Interest earned on first amount = ptr/100 = x*1*8.5/100.
Interest earned on second amount = ptr/100 = 8000-x*1*6.4 / 100
so the total interest is 8.5x/100 + 6.4*(8000-x)/100
(8.5x+51200-6.4x) = 575*100
2.1x = 57500-51200 = 6300 ----> x= 6300/2.1 =3000.
So the first amount is $3000. and second amount is $5000
----------------------------------------
Please dont pay if u are asked to .